how to calculate hybridisation order? give formula with suitable examples

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
As we know metals are electropositive,they form ionic bond .When we talk about hybridisation we generally talk about the covalent bonding.
Ex-Consider the compound MNYwhere M is the central atom and N is the atom surrounding it.
Hybridisation of such compound is given by=y+1/2(group of element M-y*valency of element N )
the number thus obtained can be put in the formula of sp1-3d1-5
for example for the number obtained as 6.hybridisation sp^3d
for charged species such as MNyK-
the hybridisation is given by =y+1/2(group of element M-y*valency of element N +k )
For positive charge put simply -k in the above formula in place of +k
I think with these formulas you can calculate the hybridisation of each and every compound
For ex-CH4
4+1/2(4-4) =4
So the hybridisation will be sp3